OpenStudy (anonymous):

Solve : If g(x) = 2(x − 4), find the value of x if g(x) = 20.

OpenStudy (lgbasallote):

just solve for x in 2(x-4) = 20 does that help?

OpenStudy (anonymous):

Uh no it dosent...

OpenStudy (anonymous):

substitute g(x) in the first equation by what it is equal to. g(x)=20

OpenStudy (anonymous):

is the answer 10?

OpenStudy (anonymous):

x-4=10 but the final answer is not 10

OpenStudy (anonymous):

hmmmmmmmmm

OpenStudy (anonymous):

12?

OpenStudy (anonymous):

\[f(x)=2(x-4)\] \[f(x)=20\] substitue f(x) with 20 in the first equation \[20=2(x-4)\] divide both sides by 2 \[\frac{ 20 }{ 2 }=x-4\] add four to each side \[10+4=x\]

OpenStudy (anonymous):

oh so it was 14?

OpenStudy (anonymous):

do you understand how we got that?

OpenStudy (anonymous):

a little

OpenStudy (anonymous):

good luck